let 1 cos A = [(m-1)(m+3)/4m] 1/2 . prove that there r 2 values for the third side one of which is m times the other.

let 1< m<3. In a triangle ABC if 2b = (m+1)a and


cos A = [(m-1)(m+3)/4m]1/2 . prove that there r 2 values for the third side one of which is m times the other.

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear

cos A = [(m-1)(m+3)/4m]1/2

calculate sinA = [1-cos2A]1/2

                    = [(m+1)(3-m)/4m]1/2

 

now sinB /b =sinA /a

sinB = (m+1)/2 sinA

         =  [(m+1)3(3-m)/16m]1/2

 

now calculate CosB = ± [1-sin2B]1/2

                                    =± [(m-1)3(m+3)/16m]1/2

now sinC =sin(180 -A-B)

                 =sin(A+B)

                 =sinA cosB + cosA sinB

 

                 =[(m-1)(m+3)(1+m)(3-m)/16m2]1/2 [(1+m) ±(m-1)]

 

 

for + value   SinC =[(m-1)(m+3)(1+m)(3-m)/16m2]1/2 [2m]

for - alue     sinC =[(m-1)(m+3)(1+m)(3-m)/16m2]1/2 [(2]

 

clearly one value is m times the other

Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin